OpenStudy (happpily):

Statistics problem! Can some please teach me how to solve this? Victoria has a bag containing 64 jellybeans. The jellybeans are red, white, black, and yellow. The probability of pulling a white jellybean from the bag is 25%. The probability of pulling a red is 1 out of 16. There are the same numbers of yello jellybeans as black. If a black jellybean is pulled out and eaten, what is the probability that the next jellybean to be pulled out will be Black?

OpenStudy (faiqraees):

\[\large \color{black}{\begin{align} \begin{array}{|c|c|c|} \hline Bean & Red & Black & White & Yellow \ \\ \hline Probability &1/16 & & 0.25 & \\ \hline \end{array} \end{align}}\] Can you complete the table?

OpenStudy (happpily):

B = 22 Y= 22

OpenStudy (faiqraees):

@Happpily Well I was looking for PROBABILITY but that will aslo work

OpenStudy (faiqraees):

\[ \large \color{black}{\begin{align} \begin{array}{|c|c|c|} \hline Bean & Red & Black & White & Yellow \ \\ \hline Probability &1/16 &22/64 & 0.25 & 22/64\\ \hline \end{array} \end{align}} \]

OpenStudy (happpily):

Do I have to divide that?

OpenStudy (faiqraees):

What will be the probability to pick a Black bean after one pick?

OpenStudy (happpily):

21/63

OpenStudy (faiqraees):

very good P(Black and Black) = P(Black) x P(Black) = 22/64 x 21/63

OpenStudy (happpily):

Do I cross multiply?

OpenStudy (faiqraees):

No this is just fraction multiplication \[\large\rm \frac{22}{64}*\frac{21}{63} \]

OpenStudy (happpily):

462/4,032

OpenStudy (faiqraees):

Simplify

OpenStudy (happpily):

8.727

OpenStudy (faiqraees):

No

OpenStudy (happpily):

11/96

OpenStudy (faiqraees):

Correct

OpenStudy (happpily):

So thats the answer?

OpenStudy (faiqraees):

Yes

OpenStudy (faiqraees):

No I think it isn't that complicated they are just asking the probability of the second black jellybean which is 21/63 = 1/3 Which is the answer

OpenStudy (happpily):

Yeah that's all they're asking. But what does the 11/96 have to do with it?

OpenStudy (happpily):

I'm a bit confused now lol

OpenStudy (faiqraees):

leave that

OpenStudy (happpily):

Just forget about it? How did you get 1/3 from 21/63?

OpenStudy (faiqraees):

21÷7/63÷7 = 1/3 And can you rephrase your second question

OpenStudy (happpily):

Nevermind. Thank you so much for your help. If it's not too much to ask can you rerun the steps for me to find the probability of the black jellybean after it is eaten. I got kind of confused after 11/96